projectile motion

A projectile of mass m1 = 0.1 kg makes an inelasti
Answer #1

media%2F0a7%2F0a7b0cc6-c926-43d2-8d13-b7

Answer question
Answer #2

height=0.5(1-cos(37))=0.1006m

(m1+m2)*g*0.1006=0.5*m1*v^2

v=4.657